The sum of two consecutive even integers is ten less than twice their product. Find their integers using factoring.
Answer by nerdybill(5411)

The sum of two consecutive even integers is ten less than twice their product. Find their integers using factoring.
.
Let x = 1st consecutive even integers
then
x+2 = 2nd consecutive even integers
.
x + x+2 = 2(x)(x+2) -10
2x+2 = 2x(x+2) -10
2x+2 = 4x^2+4x -10
2 = 4x^2+2x-10
0 = 4x^2+2x-12
0 = 2x^2+x-6
.
0 = (2x-3)(x+2)
.
Setting each to zero:
2x-3 = 0
2x = 3
x = 3/2
.
x+2 = 0
x = -2
.
Solution: x = {-2, 3/2}